# PhotoDoodle: Learning Artistic Image Editing from Few-Shot Pairwise Data

### v1

@@ 1 Introduction
The rise of diffusion models has started a new chapter for image creation and control. Using the pretrain-finetune approach, the community has achieved remarkable progress in customized image generation [25, 10, 12], with applications spanning identity preservation [18], artistic stylization [30, 45, 28, 2], and subject coherence [16, 25, 10, 12, 26, 14, 48]. However, while customized image generation thrives, a critical gap persists in customized image editing. This problem contrasts sharply with the growing need for precise image editing tools, as current methods mainly focus on creating new content rather than making smart edits based on context. This less explored area of custom image editing presents both technical challenges and chances to create new important applications.
Photo Doodling represents a long-recognized yet historically challenge in customized image editing. This task involves artists enhancing background photographs through the strategic integration of decorative elements and stylized modifications to achieve personalized or artistic effects. Typical techniques include localized stylization, decorative linework, novel object insertion, and ornamental augmentation of existing subjects. Artists and designers exhibit distinctive personal styles and strategic approaches during the photo doodling creation process, which holds significant value. However, the absence of automated tools forces manual workflows requiring hour-level time investments per image, severely constraining both productivity and the feasibility of generating large-scale paired training data.
Automating PhotoDoodle workflows is intriguing, yet it faces three interconnected challenges. First, achieving harmonious integration demands that generated decorations align not only spatially (e.g., matching perspective) but also semantically with the background, while accommodating non-photorealistic styles such as cartoon or flat illustration. Second, maintaining strict background consistency requires rigorously preventing unintended alterations to the original content, including color shifts or texture distortions. Third, to achieve efficient style acquisition, it is necessary to extract an artist’s unique editing patterns from a few photo-doodle pairs, which is technically challenging. These compounded challenges cause existing methods to be incompetent in addressing the problem comprehensively.
@@
The limitations of prevailing image editing paradigms further exacerbate these difficulties. Global editing methods (e.g., Prompt-to-Prompt[11], InstructP2P[5]), while effective for consistent style transfer, inadvertently distort background content during localized modifications. Inpainting-based approaches, though capable of preserving non-mask regions through localized denoising, impose impractical demands for pixel-perfect user-defined masks, fundamentally conflicting with PhotoDoodle’s need for flexible creative workflows. To bridge this gap, we introduce an instruction-guided framework that replaces mask dependency with semantic-level control, enabling precise and style-conscious decorative generation while ensuring background consistency.
To address these challenges and bridge the gaps in existing solutions, we propose PhotoDoodle, a framework designed to learn artistic image editing from few-shot examples. Building upon a pre-trained Diffusion Transformer (DiT), our method conditions on background images to generate photo doodles through a two-stage training pipeline. First, we adapt a well-trained text-to-image model into a versatile image editor by fine-tuning it on large-scale editing datasets, thereby establishing the foundational editing model OmniEditor. Then, we train an EditLoRA module using artist-curated before-after pairwise data to learn nuanced editing styles, while introducing a Positional Encoding (PE) Cloning strategy to enforce strict background consistency through implicit feature alignment.
Our design is guided by two main insights. First, maintaining clean latent conditioning—ensuring that input image tokens are noise-free—is essential for keeping the output coherent with the background by minimizing divergence in the latent space. Second, implicit alignment through position encoding cloning enables the seamless integration of background context without adding extra trainable parameters, thus preserving the full generative capacity of the base T2I model. These insights allow our framework to strike a balance between artistic flexibility and strict consistency.
@@
  • We propose a customizable image editing framework that learns and transfers artistic styles from few photo-doodle pairs.
  • By extending conditional mechanisms and implementing a PE cloning strategy for implicit alignment, we achieve high-quality and highly consistent image editing built upon Diffusion Transformers (DiTs), while utilizing EditLoRA to efficiently learn customized editing operations.
  • We propose the first dedicated photo-doodle dataset containing 300+ high-quality pairs across 6 artistic styles, establishing a benchmark for reproducible research.
@@ 2.1 Diffusion Model and Conditional Generation
Recent advances in diffusion models have significantly advanced the state of text-conditioned image synthesis, achieving remarkable equilibrium between generation diversity and visual fidelity. Pioneering works such as GLIDE [20], hierarchical text-to-image models [23], and photorealistic synthesis frameworks [27] have systematically addressed key challenges in cross-modal generation tasks. The emergence of Stable Diffusion [24], which implements a Latent Diffusion Model with text-conditioned UNet architecture, has established new benchmarks in text-to-image generation and inspired subsequent innovations including SDXL [22], GLiGEN [17], and Ranni [9]. To enhance domain-specific adaptation, parameter-efficient fine-tuning techniques like Low-Rank Adaptation (LoRA) [12] and DreamBooth [25] have demonstrated effective customization of pre-trained models. Concurrently, research efforts have focused on precise control over pictorial concepts through multi-concept customization [kumari2023multi], image prompt integration [39], and identity-preserving generation [36]. The introduction of ControlNet [43] further extended controllable generation capabilities to spatial constraints and depth information, with subsequent optimizations improving inference efficiency [46, 47]. Some work [35, 7, 32, 33] also focuses on the security issues of customized generation. Building upon these foundations, our work investigates the novel application of pre-trained diffusion transformers for generating artistic photo-doodles. Unlike previous approaches focusing on photorealism or explicit control modalities, we explore the model’s potential in capturing freehand artistic expression while maintaining structural coherence.
Figure 2: The overall architecture and training prodigim of photodoodle. The ominiEditor and EditLora all follow the lora training prodigm. We use a high rank lora for pre-training the OmniEditor on a large-scale dataset for general-purpose editing and text-following capabilities, and a low rank lora for fine-tuning EditLoRA on a small set of paired stylized images to capture individual artists’ specific styles and strategies for efficient customization. We encode the original image into a condition token and concatenate it with a noised latent token, controlling the generation outcome through MMAttention.
@@ 2.2 Text-guilded Image Editing
Recent advances in text-guided image editing have established this field as a critical research frontier in visual content manipulation, with current methodologies generally classified into three paradigms: global description-guided, local description-guided, and instruction-guided editing. Global description-guided approaches (e.g., Prompt2Prompt [11], Imagic [15], EditWorld [38] ), achieve fine-grained manipulation through cross-modal alignment between textual descriptions and image regions, yet demand meticulous text specification for target attributes. Local description-guided methods such as Blended Diffusion [3] and Imagen Editor [37] enable pixel-level control via explicit mask annotations and regional text prompts, though their practical application is constrained by the requirement for precise spatial specifications, particularly in complex editing scenarios like object removal. The emerging instruction-guided paradigm, exemplified by InstructPix2Pix [5] and HIVE [44], represents a paradigm shift through its natural language interface that accepts editing commands (e.g., ”change the scene to spring”). This approach eliminates the dependency on detailed textual descriptions or precise mask annotations, significantly enhancing user accessibility.
@@ 3.4 EditLoRA
Figure 4: Compared to baseline methods, PhotoDoodle demonstrates superior instruction following, image consistency, and editing effectiveness.
@@ 4.2 Generation Results
Fig. 3 displays the image editing results of PhotoDoodle, which excels in text following due to training on a large dataset of before-after pairs. The generated doodles blend well with the original images. When trained on a limited dataset of artist-paired data using EditLoRA, PhotoDoodle consistently produces artistic doodles while preserving image consistency and avoiding unwanted changes. Notably, our method maintains stable performance and a high success rate, making it suitable for production use and reducing the need for selective sampling.
@@ 4.2.1 Qualitative Evaluation
Figure 5: Ablation study results. Without OmniEditor pre-training, training EditLoRA directly reduces the harmony between sketches and photos and weakens text-following capabilities; removing Position Encoding Cloning decreases output consistency and introduces unwanted background changes; using only pre-trained OmniEditor without EditLoRA significantly reduces the degree of stylization.

### v2

# PhotoDoodle: Learning Artistic Image Editing from Few-Shot Examples

@@ 1 Introduction
The rise of diffusion models has started a new chapter for image creation and control. Using the pretrain-finetune approach, the community has achieved remarkable progress in customized image generation [28, 10, 14], with applications spanning identity preservation [21], artistic stylization [33, 48, 31, 2], and subject coherence [18, 28, 10, 14, 29, 16, 52]. These advancements have fueled applications ranging from digital art creation to commercial design.
Despite these successes, there remains a critical gap between customized image generation and image editing. Current methods primarily focus on content creation, while intelligent context-aware editing, particularly for artistic enhancement, remains underexplored. This imbalance stands in contrast to the growing demand for precision image editing tools, with current approaches focusing on content generation rather than context-aware editing. As a result, this underexplored frontier of customized image editing is both a technical challenge and a key to unlocking transformative applications.
PhotoDoodling, as a paradigmatic challenge in customized image editing, demands artists to enhance background photographs through strategic integration of decorative elements (e.g., stylized linework, ornamental patterns) and context-aware modifications for personalized aesthetics. Conventional workflows involve artistic techniques, such as: (1) local stylization, (2) decorative contour rendering, (3) semantic-aware object insertion, and (4) ornamental augmentation. While these processes showcase distinctive artistic signatures and strategic design logic, their manual execution incurs prohibitive time costs, fundamentally constraining both production scalability and the curation of large-scale paired training datasets required for data-driven approaches. However, automating these workflows introduces three interlocked technical barriers: First, harmonious integration demands generated decorations to simultaneously satisfy perspective alignment and semantic coherence with background contexts. Second, strict background preservation requires mechanisms to prevent unintended changes, such as color distribution shifting and texture pattern alternation. Third, efficient style distillation must extract artists’ unique editing patterns from sparse pairwise examples (30-50 image pairs). These compounded challenges cause existing methods to be incompetent in addressing the problem in a comprehensive way.
@@
Prevailing image editing paradigms can hardly deal with these challenges altogether. Global editing methods (e.g., Prompt-to-Prompt[13], InstructP2P[5]), while effective for consistent style transfer, inadvertently distort background content during localized modifications. Inpainting-based approaches[45, 53], though capable of preserving unmasked regions through localized editing, impose impractical demands for pixel-perfect user-defined masks, fundamentally conflicting with the need for automatic PhotoDoodling. To bridge this gap, we introduce an instruction-guided image editing framework that discarded mask dependency, enabling precise and style-conscious decorative generation while ensuring background consistency.
Our proposed framework, PhotoDoodle, presents a few-shot artistic image editing framwork built upon Diffusion Transformers (DiT), featuring a dual-stage training architecture. In the first phase, we evolve a pre-trained text-to-image DiT model into a universal image editor (OmniEditor) through two key innovations: (1) a Positional Encoding (PE) Cloning mechanism that preserves spatial fidelity by providing coordinate-aware hints, and (2) a noise-free conditioning paradigm that offers non-distorted information of the source image. This foundational stage is trained on 3.5M image editing pairs[11], establishing robust general editing capabilities. The second phase introduces an EditLoRA module that distills artist-specific editing patterns from merely 30-50 exemplar pairs through low-rank adaptation(LoRA), enabling efficient style customization while maintaining base model’s capability. This co-designed architecture ensures a balance between artistic flexibility and strict consistency.
@@
  • First framework for artistic photo-doodling: A DiT-based architecture to enable few-shot learning of style-specific editing operations while preserving background integrity.
  • We propose a noise-free conditioning paradigm with positional encoding cloning for implicit feature alignment, enabling high-fidelity image editing through EditLoRA-enhanced Diffusion Transformers (DiTs) that efficiently learn customized operations while maintaining strict background consistency
  • We collected the first publicly available curated photo-doodle dataset comprising 300+ high-quality pairs across 6 artistic styles, establishing a benchmark for reproducible research.
@@ 2.1 Diffusion Model and Conditional Generation
Recent advances in diffusion models have significantly advanced the state of text-conditioned image synthesis, achieving remarkable equilibrium between generation diversity and visual fidelity. Pioneering works such as GLIDE [23], hierarchical text-to-image models [26], and photorealistic synthesis frameworks [30] have systematically addressed key challenges in cross-modal generation tasks. The emergence of Stable Diffusion [27], which implements a Latent Diffusion Model with text-conditioned UNet architecture, has established new benchmarks in text-to-image generation and inspired subsequent innovations including SDXL [25], GLiGEN [20], and Ranni [9]. To enhance domain-specific adaptation, parameter-efficient fine-tuning techniques like Low-Rank Adaptation (LoRA) [14] and DreamBooth [28] have demonstrated effective customization of pre-trained models. Concurrently, research efforts have focused on precise control over pictorial concepts through multi-concept customization [kumari2023multi], image prompt integration [ye2023ip,ssr, 49], and identity-preserving generation [39]. The introduction of ControlNet [46] further extended controllable generation capabilities to spatial constraints and depth information, with subsequent extended applications on various scenarios[51, 40, 50, 12]. Some work [38, 7, 35, 36] also focuses on the security issues of customized generation. Building upon these foundations, our work investigates the novel application of pre-trained diffusion transformers for generating artistic photo-doodles. Unlike previous approaches focusing on photorealism or explicit control modalities, we explore the model’s potential in capturing freehand artistic expression while maintaining structural coherence.
Figure 2: The overall architecture and training prodigim of photodoodle. The ominiEditor and EditLora all follow the lora training prodigm. We use a high rank lora for pre-training the OmniEditor on a large-scale dataset for general-purpose editing and text-following capabilities, and a low rank lora for fine-tuning EditLoRA on a small set of paired stylized images to capture individual artists’ specific styles and strategies for efficient customization. We encode the source image into a condition token and concatenate it with a noised latent token, controlling the generation outcome through MMAttention.
@@ 2.2 Text-guilded Image Editing
Recent advances in text-guided image editing have established this field as a critical research frontier in visual content manipulation, with current methodologies generally classified into three paradigms: global description-guided, local description-guided, and instruction-guided editing. Global description-guided approaches (e.g., Prompt2Prompt [13], Imagic [17], EditWorld [42], ZONE [19] ), achieve fine-grained manipulation through cross-modal alignment between textual descriptions and image regions, yet demand meticulous text specification for target attributes. Local description-guided methods such as Blended Diffusion [3] and Imagen Editor [41] enable pixel-level control via explicit mask annotations and regional text prompts, though their practical application is constrained by the requirement for precise spatial specifications, particularly in complex editing scenarios like object removal. The emerging instruction-guided paradigm, exemplified by InstructPix2Pix [5] and HIVE [47], represents a paradigm shift through its natural language interface that accepts editing commands (e.g., ”change the scene to spring”). This approach eliminates the dependency on detailed textual descriptions or precise mask annotations, significantly enhancing user accessibility.
@@ 3.4 EditLoRA
Figure 4: Compared to baselines, PhotoDoodle demonstrates superior instruction following, image consistency, and editing effectiveness.
@@ 4.2 Generation Results
Fig. 3 and Fig. 7 displays the image editing results of PhotoDoodle, which excels in text following due to training on a large dataset of before-after pairs. The generated doodles blend well with the original images. When trained on a limited dataset of artist-paired data using EditLoRA, PhotoDoodle consistently produces artistic doodles while preserving image consistency and avoiding unwanted changes. Notably, our method maintains stable performance and a high success rate, making it suitable for production use and reducing the need for selective sampling.
@@ 4.4 User Study
Figure 7: More photo doodling results: one adds lines to a photo of clouds, imagining them as animals; the other converts the photo into a monochrome version and decorates it with color blocks.